OpenStudy (gschibby):

Hi! I need some help with Laplace transformations :) It's pretty basic, I just need to know how to transform t-3 :P

OpenStudy (gschibby):

\[L\left\{ t-3\right\}=?\]

OpenStudy (gschibby):

Do I separate it like this:\[L \left\{ t-3 \right\}=L \left\{ t \right\}-L \left\{ 3 \right\}\]?

OpenStudy (gschibby):

or\[\lim_{T \rightarrow \infty}\int\limits_{0}^{T}(t-3)e ^{-st}dt\]

ganeshie8 (ganeshie8):

just plug that into ithe integral and evaluate ?

OpenStudy (gschibby):

Hm, ok, I'll try :P

ganeshie8 (ganeshie8):

or you could use \[\large 1\leadsto \dfrac{1}{s}\]

ganeshie8 (ganeshie8):

and the fact that the transform is linear

ganeshie8 (ganeshie8):

L(3) = 3*L(1) = 3/s

OpenStudy (gschibby):

So\[L \left\{ t-3 \right\}=L \left\{ t \right\}-L \left\{ 3 \right\}=L \left\{ t \right\}-3L \left\{ 1 \right\}\]\[=\frac{ 1 }{ s ^{2} }-\frac{ 3 }{ s }\] ?

ganeshie8 (ganeshie8):

looks good !

OpenStudy (gschibby):

Thanks!^^

ganeshie8 (ganeshie8):

np :)